OpenStudy (anonymous):

PLEASE ANSWER THE FIRST PART IF YOU CAN ):!Suppose you play a game of chance where you roll a pair of dice. You pay a nonrefundable $2 to play the game. If the number of dots on both dice is even, you win $6. Otherwise you win nothing. 1) what is the probability of winning. 2) create a probability distribution table for X where is X is the random variable.

OpenStudy (dumbcow):

the dice must both be even, or show a 2,4,or 6 this leaves only 3 options for each die ---> 3*3 = 9 possible winning combinations There are a total of 6*6=36 combinations probability = 9/36 = 1/4

OpenStudy (anonymous):

thank you soo much!

OpenStudy (dumbcow):

yw i wouldnt play this game .... on avg you will lose 50 cents per game :)

OpenStudy (anonymous):

lmaooo haha yea i dont think anyone would XD

OpenStudy (dumbcow):

yet millions play the lottery ...go figure haha

OpenStudy (anonymous):

false hope is good sometimes xD